### v2

@@ A.2 $D^{*}\bar{B}\to D^{*}\bar{B}$
Figure 13: LO contact, NLO contact and NLO TPE diagrams of the process $D^{*}\bar{B}\to D^{*}\bar{B}$.
For the $D^{*}(p_{1})\bar{B}(p_{2})\to D^{*}(p_{3})\bar{B}(p_{4})$ process, the scattering amplitude of LO contact diagram is
@@
with isospin $I=0$, where $\epsilon_{1}$ and $\epsilon^{*}_{4}$ are the polarization vectors of the iniatial $D$ and final $\bar{B}^{*}$, respectively. Next, we consider the NLO interactions in the $D^{*}(p_{1})\bar{B}(p_{2})\to D^{*}(p_{3})\bar{B}^{(}p_{4})$ process. The corresponding diagrams are illustrated in Fig. 13. The one-loop corrections to the contact term are:
@@
×$\times(\epsilon_{1}\cdot\epsilon^{*}_{3}),$ (150)
where coefficients appearing in the above contact amplitudes are shown in Table 15. The TPE amplitudes at NLO are given by:
$\mathcal{M}_{c5.1}^{(2)}=$ $~{}4\frac{1}{f^{4}}[A_{1}(q_{0}^{2}J_{0}^{F}+J_{22}^{F})+A_{15}(q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})-A_{51}(q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})+A_{5}(q_{0}^{2}J_{0}^{F}+2q_{0}^{2}J_{11}^{F}$
@@
$\mathcal{M}_{c5.2}^{(2)}=$ $-i4\frac{g_{1}^{2}}{f^{4}}\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}+(1-d)q_{0}J_{31}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}+(1-d)J_{34}^{T}]-A_{5}[q_{0}\vec{q}^{2}J_{11}^{T}$
$+(1-d)q_{0}(J_{21}^{T}+J_{31}^{T})+2q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}+(1-d)J_{34}^{T}]\}(\epsilon_{1}\cdot\epsilon^{*}_{3}),$ (152)
$\mathcal{M}_{c5.3}^{(2)}=$ $-i4\frac{g_{1}^{2}}{f^{4}}(d-3)[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})$
$-i4\frac{g_{1}^{2}}{f^{4}}(d-3)\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{S}+\vec{q}^{2}J_{24}^{S}+(2-d)q_{0}J_{31}^{S}+q_{0}\vec{q}^{2}J_{32}^{S}+\vec{q}^{2}J_{33}^{S}+(2-d)J_{34}^{S}]-A_{5}[q_{0}\vec{q}^{2}J_{11}^{S}$
$+(2-d)q_{0}(J_{21}^{S}+J_{31}^{S})+2q_{0}\vec{q}^{2}J_{22}^{S}+\vec{q}^{2}J_{24}^{S}+q_{0}\vec{q}^{2}J_{32}^{S}+\vec{q}^{2}J_{33}^{S}+(2-d)J_{34}^{S}]\}(\epsilon_{1}\cdot\epsilon^{*}_{3}),$ (153)
$\mathcal{M}_{c5.4}^{(2)}=$ $~{}i4\frac{g_{1}^{2}}{f^{4}}\{[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})$
$+[A_{1}(q_{0}J_{31}^{S}+J_{34}^{S})-A_{5}(q_{0}J_{21}^{S}+q_{0}J_{31}^{S}+J_{34}^{S})](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (154)
$\mathcal{M}_{c5.5}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)J_{31}^{B}-2\vec{q}^{2}J_{32}^{B}+(d+3)J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})$
$+[-\vec{q}^{2}J_{21}^{B}+\vec{q}^{4}J_{22}^{B}-(2d+1)\vec{q}^{2}(J_{31}^{B}+J_{42}^{B})+2\vec{q}^{4}J_{32}^{B}+(d+1)(d-2)J_{41}^{B}+\vec{q}^{4}J_{43}^{B}](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (155)
$\mathcal{M}_{c5.6}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)(J_{31}^{B}+J_{42}^{B})-2\vec{q}^{2}J_{32}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})-[\vec{q}^{2}(J_{31}^{B}$
$+J_{42}^{B})-(1+d)J_{41}^{B}](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (156)
$\mathcal{M}_{c5.7}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)J_{31}^{R}-2\vec{q}^{2}J_{32}^{B}+3J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})+[-\vec{q}^{2}J_{21}^{R}$
$+\vec{q}^{4}J_{22}^{R}-(2d+1)\vec{q}^{2}(J_{31}^{R}+J_{42}^{R})+2\vec{q}^{4}J_{32}^{R}+(d+1)(d-2)J_{41}^{R}+\vec{q}^{4}J_{43}^{R}](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (157)
$\mathcal{M}_{c5.8}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)(J_{31}^{R}+J_{42}^{R})-2\vec{q}^{2}J_{32}^{R}-\vec{q}^{2}J_{43}^{R}](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})-[\vec{q}^{2}J_{31}^{R}$
$-(1+d)J_{41}^{R}+\vec{q}^{2}J_{42}^{R}](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (158)
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{a5.1}$ 00 00 00 00
$A_{a5.2}$ 00 00 $\delta_{1}$ $\delta_{1}$
$A_{a5.3}$ 00 00 $-\delta_{2}$ $-\delta_{2}$
$A_{a5.4}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $-\delta_{2}$ 00
$A_{a5.5}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $-\delta_{2}$
$A_{a5.6}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $-\delta_{2}$ $\delta_{1}$
$A_{a5.7}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{1}$ $-\delta_{2}$
$A_{a5.8}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ 00 $-\delta_{2}$
$A_{a5.(9+10)}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $\delta_{1}$ 00
@@
      $A_{c5.1}$       $\frac{1}{16}$       $\frac{1}{16}$       $-\frac{1}{16}$       $-\frac{3}{16}$       $-\frac{3}{16}$       $\frac{3}{16}$       00       00
      $A_{c5.2}$       $\frac{i}{8}$       $\frac{-i}{8}$       00       $-\frac{3i}{8}$       $\frac{3i}{8}$       00       $-\delta_{2}$       00
      $A_{c5.3}$       $\frac{i}{8}$       $-\frac{i}{8}$       00       $-\frac{3i}{8}$       $\frac{3i}{8}$       00       00       00
      $A_{c5.4}$       $\frac{i}{8}$       $-\frac{i}{8}$       00       $-\frac{3i}{8}$       $\frac{3i}{8}$       00       $\delta_{1}$       00
      $A_{c5.5}$       $\frac{1}{16}$       00       00       $\frac{9}{16}$       00       00       00       $-\delta_{2}$
      $A_{c5.6}$       $\frac{1}{16}$       00       00       $\frac{9}{16}$       00       00       $\delta_{1}$       $-\delta_{2}$
      $A_{c5.7}$       $\frac{5}{16}$       00       00       $-\frac{3}{16}$       00       00       00       $-\delta_{2}$
      $A_{c5.8}$       $\frac{5}{16}$       00       00       $-\frac{3}{16}$       00       00       $\delta_{1}$       $-\delta_{2}$
@@ A.3 $D\bar{B}^{*}\to D^{*}\bar{B}$
Figure 14: LO contact, NLO contact, LO OPE, NLO OPE and NLO TPE diagrams of the process $D\bar{B}^{*}\to D^{*}\bar{B}$.
Now we focus on the process $D(p_{1})\bar{B}^{*}(p_{2})\to D^{*}(p_{3})\bar{B}(p_{4})$, the corresponding diagrams are illustrated in Fig. 14. The scattering amplitudes of LO diagrams are given by
@@
The one-loop corrections to the contact amplitude are:
@@
×$\times(\epsilon_{2}\cdot\epsilon^{*}_{3}),$ (171)
@@
×$\times(\epsilon_{2}\cdot\epsilon^{*}_{3}),$ (173)
@@
From Fig. 14, we can obtain the scattering amplitudes of NLO OPE diagrams:
@@
×$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (181)
@@
×$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (182)
@@
×$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (184)
@@
×$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (186)
@@
×$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (187)
@@
×$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (188)
@@
×$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3}.$ (189)
The TPE amplitudes are given by:
@@
In Tables 17-18, we list the coefficients of the contact amplitudes and OPE amplitudes at NLO, respectively. For the TPE terms, the coefficients are
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{a6.1(2)}$ 00 00 $0(\delta_{2})$ $-\delta_{2}$
$A_{a6.3(4)}$ 00 00 $-\delta_{1}$ $0(\delta_{1})$
$A_{a6.5}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $-\delta_{2}$
$A_{a6.6}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta_{2}$ $-\delta_{1}$
$A_{a6.7}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $-\delta_{1}$
$A_{a6.8}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta_{1}$ $-\delta_{2}$
$A_{a6.(9+10)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{1}$
$A_{a6.11}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $-\delta_{1}$ 00
$A_{a6.(12+13)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{2}$
$A_{a6.14}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $-\delta_{2}$ 00
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{b6.1(2)}$ $-\frac{1}{16}$ $\frac{3}{16}$ $0(\delta_{2})$ $-\delta_{2}$
$A_{b6.3(4)}$ $-\frac{1}{16}$ $\frac{3}{16}$ $-\delta_{1}$ $0(\delta_{1})$
$A_{b6.5(6)}$ $-\frac{1}{12}$ $\frac{1}{4}$ 00 00
$A_{b6.7}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b6.(8+9)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 $\delta_{1}$
$A_{b6.10}$ $\frac{1}{4}$ $-\frac{3}{4}$ $-\delta_{1}$ 00
$A_{b6.(11+12)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 $\delta_{2}$
$A_{b6.13}$ $\frac{1}{4}$ $-\frac{3}{4}$ $-\delta_{2}$ 00
$A_{b6.14(15)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b6.16(17)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00

### v3

@@ A.2 $D^{*}\bar{B}\to D^{*}\bar{B}$
Figure 13: LO contact ($a5$), NLO contact (a5.1a5.10a5.1-a5.10) and NLO TPE (c5.1c5.8c5.1-c5.8) diagrams of the process $D^{*}\bar{B}\to D^{*}\bar{B}$.
For the $D^{*}(p_{1})\bar{B}(p_{2})\to D^{*}(p_{3})\bar{B}(p_{4})$ process, the scattering amplitude of the LO contact diagram is
@@
with isospin $I=0$, where $\epsilon_{1}$ and $\epsilon^{*}_{4}$ are the polarization vectors of the initial $D$ and final $\bar{B}^{*}$, respectively. Next, we consider the NLO interactions in the $D^{*}(p_{1})\bar{B}(p_{2})\to D^{*}(p_{3})\bar{B}^{(}p_{4})$ process. The corresponding diagrams are illustrated in Fig. 13. The one-loop corrections to the contact term are
@@
$\times(\epsilon_{1}\cdot\epsilon^{*}_{3}),$ (150)
where coefficients appearing in the above contact amplitudes are shown in Table 15. The TPE amplitudes at NLO are given by
$\mathcal{M}_{c5.1}^{(2)}=$ $~4\frac{1}{f^{4}}[A_{1}(q_{0}^{2}J_{0}^{F}+J_{22}^{F})+A_{15}(q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})$
$-A_{51}(q_{0}^{2}J_{11}^{F}+q_{0}^{2}J_{21}^{F}+J_{22}^{F})+A_{5}(q_{0}^{2}J_{0}^{F}+2q_{0}^{2}J_{11}^{F}$
@@
$\mathcal{M}_{c5.2}^{(2)}=$ $-i4\frac{g_{1}^{2}}{f^{4}}\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}+(1-d)q_{0}J_{31}^{T}$
$+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}+(1-d)J_{34}^{T}]$
$-A_{5}[q_{0}\vec{q}^{2}J_{11}^{T}+(1-d)q_{0}(J_{21}^{T}+J_{31}^{T})$
$+2q_{0}\vec{q}^{2}J_{22}^{T}+\vec{q}^{2}J_{24}^{T}+q_{0}\vec{q}^{2}J_{32}^{T}+\vec{q}^{2}J_{33}^{T}$
$+(1-d)J_{34}^{T}]\}(\epsilon_{1}\cdot\epsilon^{*}_{3}),$ (152)
$\mathcal{M}_{c5.3}^{(2)}=$ $-i4\frac{g_{1}^{2}}{f^{4}}(d-3)[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})$
$-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}$
$+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})$
$-i4\frac{g_{1}^{2}}{f^{4}}(d-3)\{A_{1}[q_{0}\vec{q}^{2}J_{22}^{S}+\vec{q}^{2}J_{24}^{S}$
$+(2-d)q_{0}J_{31}^{S}+q_{0}\vec{q}^{2}J_{32}^{S}+\vec{q}^{2}J_{33}^{S}+(2-d)J_{34}^{S}]$
$-A_{5}[q_{0}\vec{q}^{2}J_{11}^{S}+(2-d)q_{0}(J_{21}^{S}+J_{31}^{S})$
$+2q_{0}\vec{q}^{2}J_{22}^{S}+\vec{q}^{2}J_{24}^{S}+q_{0}\vec{q}^{2}J_{32}^{S}+\vec{q}^{2}J_{33}^{S}$
$+(2-d)J_{34}^{S}]\}(\epsilon_{1}\cdot\epsilon^{*}_{3}),$ (153)
$\mathcal{M}_{c5.4}^{(2)}=$ $~i4\frac{g_{1}^{2}}{f^{4}}\{[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})$
$-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}$
$+J_{33}^{S})](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})+[A_{1}(q_{0}J_{31}^{S}+J_{34}^{S})$
$-A_{5}(q_{0}J_{21}^{S}+q_{0}J_{31}^{S}+J_{34}^{S})](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (154)
$\mathcal{M}_{c5.5}^{(2)}=$ $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)J_{31}^{B}$
$-2\vec{q}^{2}J_{32}^{B}+(d+3)J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})$
$+[-\vec{q}^{2}J_{21}^{B}+\vec{q}^{4}J_{22}^{B}-(2d+1)\vec{q}^{2}(J_{31}^{B}+J_{42}^{B})$
$+2\vec{q}^{4}J_{32}^{B}+(d+1)(d-2)J_{41}^{B}+\vec{q}^{4}J_{43}^{B}](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (155)
$\mathcal{M}_{c5.6}^{(2)}=$ $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)(J_{31}^{B}+J_{42}^{B})$
$-2\vec{q}^{2}J_{32}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})$
$-[\vec{q}^{2}(J_{31}^{B}+J_{42}^{B})-(1+d)J_{41}^{B}](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (156)
$\mathcal{M}_{c5.7}^{(2)}=$ $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)J_{31}^{R}$
$-2\vec{q}^{2}J_{32}^{B}+3J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})$
$+[-\vec{q}^{2}J_{21}^{R}+\vec{q}^{4}J_{22}^{R}-(2d+1)\vec{q}^{2}(J_{31}^{R}+J_{42}^{R})$
$+2\vec{q}^{4}J_{32}^{R}+(d+1)(d-2)J_{41}^{R}+\vec{q}^{4}J_{43}^{R}](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (157)
$\mathcal{M}_{c5.8}^{(2)}=$ $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)(J_{31}^{R}+J_{42}^{R})$
$-2\vec{q}^{2}J_{32}^{R}-\vec{q}^{2}J_{43}^{R}](q\cdot\epsilon_{1})(q\cdot\epsilon^{*}_{3})$
$-[\vec{q}^{2}J_{31}^{R}-(1+d)J_{41}^{R}+\vec{q}^{2}J_{42}^{R}](\epsilon_{1}\cdot\epsilon^{*}_{3})\},$ (158)
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{a5.1}$ 0 0 0 0
$A_{a5.2}$ 0 0 $\delta_{1}$ $\delta_{1}$
$A_{a5.3}$ 0 0 $\delta_{2}$ $\delta_{2}$
$A_{a5.4}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{2}$ 0
$A_{a5.5}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 $\delta_{2}$
$A_{a5.6}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{2}$ $\delta_{1}$
$A_{a5.7}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{1}$ $\delta_{2}$
$A_{a5.8}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ 0 $\delta_{2}$
$A_{a5.(9+10)}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $\delta_{1}$ 0
@@
  $A_{c5.1}$   $\frac{1}{16}$   $\frac{1}{16}$   $\frac{1}{16}$   $\frac{3}{16}$   $\frac{3}{16}$   $\frac{3}{16}$   0   0
  $A_{c5.2}$   $\frac{i}{8}$   $\frac{-i}{8}$   0   $\frac{3i}{8}$   $\frac{3i}{8}$   0   $\delta_{2}$   0
  $A_{c5.3}$   $\frac{i}{8}$   $\frac{i}{8}$   0   $\frac{3i}{8}$   $\frac{3i}{8}$   0   0   0
  $A_{c5.4}$   $\frac{i}{8}$   $\frac{i}{8}$   0   $\frac{3i}{8}$   $\frac{3i}{8}$   0   $\delta_{1}$   0
  $A_{c5.5}$   $\frac{1}{16}$   0   0   $\frac{9}{16}$   0   0   0   $\delta_{2}$
  $A_{c5.6}$   $\frac{1}{16}$   0   0   $\frac{9}{16}$   0   0   $\delta_{1}$   $\delta_{2}$
  $A_{c5.7}$   $\frac{5}{16}$   0   0   $\frac{3}{16}$   0   0   0   $\delta_{2}$
  $A_{c5.8}$   $\frac{5}{16}$   0   0   $\frac{3}{16}$   0   0   $\delta_{1}$   $\delta_{2}$
@@ A.3 $D\bar{B}^{*}\to D^{*}\bar{B}$
Figure 14: LO contact ($a6$), LO OPE ($b6$), NLO contact (a6.1a6.14a6.1-a6.14), NLO OPE (b6.1b6.17b6.1-b6.17), and NLO TPE (c6.1c6.2c6.1-c6.2) diagrams of the process $D\bar{B}^{*}\to D^{*}\bar{B}$.
Now we focus on the process $D(p_{1})\bar{B}^{*}(p_{2})\to D^{*}(p_{3})\bar{B}(p_{4})$, and the corresponding diagrams are illustrated in Fig. 14. The scattering amplitudes of LO diagrams are given by
@@
The one-loop corrections to the contact amplitudes are
@@
$\times(\epsilon_{2}\cdot\epsilon^{*}_{3}),$ (171)
@@
$\times(\epsilon_{2}\cdot\epsilon^{*}_{3}),$ (173)
@@
From Fig. 14, we can obtain the scattering amplitudes of NLO OPE diagrams,
@@
$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (181)
@@
$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (182)
@@
$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (184)
@@
$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (186)
@@
$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (187)
@@
$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3},$ (188)
@@
$\times\epsilon_{2}^{\mu}\epsilon^{*\nu}_{3}.$ (189)
The TPE amplitudes are given by
@@
In Tables 17 and 18, we list the coefficients of the contact amplitudes and OPE amplitudes at NLO, respectively. For the TPE terms, the coefficients are
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{a6.1(2)}$ 0 0 $0(\delta_{2})$ $\delta_{2}$
$A_{a6.3(4)}$ 0 0 $\delta_{1}$ $0(\delta_{1})$
$A_{a6.5}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 $\delta_{2}$
$A_{a6.6}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta_{2}$ $\delta_{1}$
$A_{a6.7}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 $\delta_{1}$
$A_{a6.8}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta_{1}$ $\delta_{2}$
$A_{a6.(9+10)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{1}$
$A_{a6.11}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $\delta_{1}$ 0
$A_{a6.(12+13)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{2}$
$A_{a6.14}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $\delta_{2}$ 0
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{b6.1(2)}$ $\frac{1}{16}$ $\frac{3}{16}$ $0(\delta_{2})$ $\delta_{2}$
$A_{b6.3(4)}$ $\frac{1}{16}$ $\frac{3}{16}$ $\delta_{1}$ $0(\delta_{1})$
$A_{b6.5(6)}$ $\frac{1}{12}$ $\frac{1}{4}$ 0 0
$A_{b6.7}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b6.(8+9)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 $\delta_{1}$
$A_{b6.10}$ $\frac{1}{4}$ $\frac{3}{4}$ $\delta_{1}$ 0
$A_{b6.(11+12)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 $\delta_{2}$
$A_{b6.13}$ $\frac{1}{4}$ $\frac{3}{4}$ $\delta_{2}$ 0
$A_{b6.14(15)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b6.16(17)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0